# LLM-based speaker diarization correction: A generalizable approach

### v2

@@ 1 Introduction
Several speaker diarization tools exist[10, 11]. A popular open source method is Pyannote [12], which uses pre-trained neural networks to assign speaker labels through analysis of acoustic data. It processes audio into overlapping segments, extracting features such as mel-frequency cepstral coefficients to detect changes in speakers. Another method is the x-vector approach [13], which extracts speaker embeddings, or x-vectors, from fixed-length segments of audio using a time-delay neural network. The embeddings are clustered to assign speaker labels. End-to-end diarization models[14, 15, 16] integrate speaker recognition with transcription, allowing for simultaneous transcription and diarization. These models map audio inputs directly to speaker-attributed transcriptions. For example, the transcribe-to-diarize model [17] applies a neural network that processes audio to provide both the transcribed text and the speaker labels in one pass.
@@
Here, we fine-tune LLMs to improve diarization accuracy in ASR transcripts as a post-processing step. To fine-tune our models, we used the English Fisher corpus [31], a large audio dataset containing 1,960 hours of phone conversations, and to test them we used a held out set of the Fisher corpus and the PriMock57 dataset [32] a smaller audio dataset containing mock primary care consultations. We consider the effects of using transcripts from different ASRs, tuning separate models for each AWS, Azure, and WhisperX transcripts. In doing so, we learn that LLM-based diarization correction models are most effective when correcting transcripts produced using the same ASR as the transcripts they were fine-tuned with, limiting their generalizability. To address this constraint, we build an ensemble model, combining weights from each of the individual models, and demonstrate that such a model has the potential to serve as a reliable and ASR-agnostic diarization correction tool. This model has been published on HuggingFace and is available at https://huggingface.co/bklynhlth/WillisDiarize-v1. We additionally released a function using the methods described here for a friendlier user interface on applying diarization correction in OpenWillis [33], a Python library for digital measurement of health, that can be found at https://github.com/bklynhlth/openwillis/wiki/WillisDiarize-with-AWS-v1.0.
@@ 2.3 Data Preprocessing
Given each JSON had ASR-specific formatting, we extracted transcribed words and speaker labels into a common format. These transcripts are referred to here as the ASR transcripts. The ASR transcripts were further standardized by removing punctuation and transforming all text to lowercase. 11% of reference transcripts were missing transcripts for part of the recording, ranging from 30 seconds up to 8 minutes and 27 seconds. For these, we trimmed the ASR transcripts to correspond with their reference transcripts.
Any ASR transcripts with only one or more than two detected speakers were removed from the dataset. Some ASR transcripts, particularly those from WhisperX, had continuously repeating word sequences [38]. For example, words such as ‘and’ or phrases such as ‘that’s right’ would repeat consecutively such that the sentence would lose meaning. To minimize the effect of repeating sequences on model fine-tuning and testing, transcripts that contained more than 10 repetitions of a sequence were removed.
To create the training data, a transcript-preserving speaker transfer algorithm (TPST) [24] was used to transfer speaker labels from the reference transcripts to the ASR transcripts. The TPST algorithm accepts speaker-labeled source text and speaker-labeled target text. It aligns the two so that the source text speaker labels match the target text. The resulting transcripts are referred to here as the oracle transcripts. Oracle transcripts have the same text as the ASR transcripts but the speaker labels of the reference transcripts. They allow for comparison of diarization accuracy between an ASR transcripts and the ground-truth speaker labeling in the reference transcripts by removing effects for differences in the transcription itself.
@@ 2.4.2 ASR-specific models
Each of the three ASR-specific models were fine-tuned for 2 epochs with a batch size of 6 and 5 gradient accumulation steps. We used Quantized Low-Ranked Adaption of Language Models [40] and Flash attention [41] for efficient fine-tuning, balancing computational resources with model performance improvement. The three resulting fine-tuned models are referred to here as the AWS model, Azure model, and WhisperX model, or collectively as the ASR-specific models.
@@ 2.5 Completion parser
A completion parser was developed to post-process the model output. It extracts only the relevant part from the model output, removing pre- and post-fixes if present, extracts speaker labels from the model output, and transfers those labels back to input text. This completion parser was integrated as part of the model’s methods to ensure no word substitutions or modifications were made through use of the model.
@@ 2.6.1 Measurement of accuracy
Diarization accuracy was measured using two established metrics: delta concatenated minimum-permutation word error rate (deltaCP) [44] and delta speaker-attributed word error rate (deltaSA) [45]. For each, word error rate (WER) first needed to be calculated. WER was measured as the percentage of transcription errors divided by the total number of words. When measuring accuracy in this manuscript, deltaCP and deltaSA were calculated with the reference transcripts serving as ground truth.
To calculate deltaCP, concatenated minimum-permutation word error rate (cpWER) was first calculated. To do this, each speaker’s words were separated into their own transcripts. For each possible permutation of the speaker’s words, the WER was calculated against the reference speaker transcript. The smallest observed WER for each of the two speakers was averaged. cpWER is considered the minimum possible WER among all permutations of the two speakers. deltaCP is simply cpWER subtracted by the original WER. The difference allows us to analyze errors introduced by speaker labeling, independent of the original WER.
To calculate deltaSA, speaker-attributed word error rate (SA-WER) was first calculated. To do this, each speaker’s words were separated into their own transcripts. WER was calculated against the reference speaker-transcripts. The WER of each of the two speakers was averaged. deltaSA is simply SA-WER subtracted by the WER.
@@ 2.6.2 Assessing performance
Model performance was assessed by calculating deltaCP and deltaSA before and after diarization correction. Six different models for diarization correction were used:
@@ 3.2 Word error rate
We looked at WER across the three ASRs. On the Fisher test set, AWS had a WER of 22.04%, Azure had a WER of 16.99%, and WhisperX had a WER of 22.39%. On the PriMock57 dataset, AWS had a WER of 12.43%, Azure had a WER of 14.28% and WhisperX had a WER of 18.69%. The PriMock57 dataset had considerably better audio quality, given the mock sessions were carefully conducted in an isolated environment, which explains the improved transcription accuracy of the Fisher test set. Azure’s Speech to Text model performed best in terms of transcription accuracy in the Fisher test set, while AWS’s Transcribe performed the best in the PriMock57 dataset. WhisperX performed the worst in both datasets.
@@ 3.3 Baseline diarization accuracy
Diarization accuracy was compared between ASR transcripts and corresponding reference transcripts without any diarization correction. In the Fisher dataset, AWS performed best with a deltaCP of 0.93% and deltaSA of 2.5%. Azure was next with a deltaCP of 1.91% and delta SA of 3.06%. WhisperX performed worst, with a deltaCP of 4.46% and deltaSA of 5.77%. On the PriMock57 dataset Azure performed the best with a deltaCP of 0.82% and deltaSA of 1.97%. AWS had a deltaCP of 2.18% and a deltaSA of 3.56%. WhisperX performed worst on the PriMock57 dataset as well with a deltaCP of 4.96% and a deltaSA of 6.56%. The difference in diarization performance between AWS’s Transcribe and Azure’s Speech to Text is indicative of the fact that different ASRs may work better for different use cases.
@@ 3.5 Fine-tuned model performance
The performance of the ASR-specific models and the ensemble model in improving diarization accuracy was measured against baseline in the Fisher test set. The results are shown in Table 2.
@@
Fine-tuned models markedly improved diarization accuracy compared to baseline in most cases.
@@
The ensemble model achieved the best overall performance. In the AWS transcripts, it had marginally worse performance compared to the AWS model. However, in the Azure and WhisperX transcripts, it performed better than even the ASR-specific models for Azure and WhisperX.
@@ 4 Discussion
We acknowledge certain limitations, opening up opportunities for future work. First, a more thorough analysis of the diarization errors produced by each ASR, would clarify the need for generalized approaches and help us understand better the types of errors each LLM is more probable to correct. Second, the project was limited to transcripts in English. This restricts diarization correction to a single language. Future work could test this model on transcripts in different languages and also include additional languages in the dataset used for fine-tuning. Third, our experimentation with varied prompts was not extensive [50]. Given we were building on work from Wang et al. 2024 [24], we used the same prompts utilized in their work. Future work exploring the transferability of the model could include contextual information (e.g., this is a conversation between a doctor and a patient; this is a phone conversation between a salesperson and a customer) [51]. Finally, though our project focused on such models serving as post-processing steps, integrating multimodal data will allow for more robust diarization systems. Combining acoustic information with semantic insights from the LLM will enhance the system’s ability to label speech more effectively [28], especially in complex and noisy environments.
@@ 5 Conclusion
We investigated the use of fine-tuned LLMs to improve speaker diarization in conversational transcripts with existing speaker labels. Our findings demonstrate that LLMs can markedly improve diarization accuracy when specifically trained to do so. However, if such training is specific to one ASR tool or using a single dataset, it may constrain the model’s performance to transcripts only from that ASR tool and with similar diarization error rates. An ensemble model with knowledge across ASR tools allows for a generalizable and ASR-agnostic diarization correction tool. We believe such a model can be a useful post-processing step in applications that depend on transcription and diarization of conversations with multiple speakers. We have made these models publicly available at HuggingFace at https://huggingface.co/bklynhlth and built a Python function for easier use of the model available at https://github.com/bklynhlth/openwillis/wiki/WillisDiarize-with-AWS-v1.0.
@@ 6 Acknowledgements
Though referenced throughout the manuscript, the authors would like to acknowledge the authors of Quan et al. 2024 [24], whose work formed the foundation of this manuscript. Much of the code used during this work, including the TPST algorithm, critical to model fine-tuning and testing, was originally developed by their group and kindly shared with us for replication.

### v3

@@ 1 Introduction
Several speaker diarization methods exist [10, 11]. Traditionally, clustering‐based approaches have been widely used. One such method is the x-vector approach [12], which extracts speaker embeddings, or x-vectors, from fixed-length segments of audio using a time-delay neural network. The embeddings are clustered to assign speaker labels. Building on these foundations, earlier neural network–based methods further improved diarization performance. A popular open source method is Pyannote [13], which uses pre-trained neural networks to assign speaker labels through analysis of acoustic data. It processes audio into overlapping segments, extracting features such as mel-frequency cepstral coefficients to detect changes in speakers.
More recently, the field has shifted towards end-to-end diarization models [14, 15, 16, 17, 18, 19] which integrate speaker recognition with transcription, allowing for simultaneous transcription and diarization. These models map audio inputs directly to speaker-attributed transcriptions. Fujita et al. [14] introduced a self-attention mechanism to enhance the End-to-End Neural Diarization (EEND) model [20], showing improved performance over traditional BLSTM-based methods by directly optimizing the diarization error rate (DER), which significantly reduces errors in overlapping speech scenarios. Zhang et al. [15] proposed a fully supervised diarization system named unbounded interleaved-state recurrent neural network (UIS-RNN), which replaces traditional unsupervised clustering modules with an online generative process, leading to significant improvements in diarization error rates, particularly when trained with high-quality, time-stamped data. Kanda et al. proposed the transcribe-to-diarize model [16], which applies a neural network that processes audio to provide both the transcribed text and the speaker labels in one pass with the ability to estimate the start and end times of each word by incorporating a minimal set of learnable parameters into the model’s internal state.
@@
Here, we fine-tune LLMs to improve diarization accuracy in ASR transcripts as a post-processing step, visualized in Figure 1. To fine-tune our models, we used the English Fisher corpus [34], a large audio dataset containing 1,960 hours of phone conversations, and to test them we used a held out set of the Fisher corpus and the PriMock57 dataset [35] a smaller audio dataset containing mock primary care consultations. We consider the effects of using transcripts from different ASRs, tuning separate models for each AWS, Azure, and WhisperX transcript. Wang et al. (2024) demonstrated notable improvements in speaker diarization accuracy through fine-tuning. Our work investigates the generalizability of this method across different ASR systems. Specifically, we identify a key limitation in previous approaches: the variability in performance when applied to different ASRs. We find that LLM-based diarization correction models are most effective when correcting transcripts produced using the same ASR as the transcripts they were fine-tuned with, limiting their generalizability. To address this constraint, we build an ensemble model that combines the weights from each of the individual models, and demonstrate that such a model has the potential to serve as a reliable and ASR-agnostic diarization correction tool.
This model has been published on HuggingFace and is available at https://huggingface.co/bklynhlth/WillisDiarize-v1. We additionally released a function using the methods described here for a friendlier user interface on applying diarization correction in OpenWillis [36], a Python library for digital measurement of health, that can be found at https://github.com/bklynhlth/openwillis. An overview of the system used in a real-world application can be seen in Figure 2.
Figure 2: Overview of pipeline used to evaluate the LLM.
@@ 2.3 Data Preprocessing
Given that each JSON file had ASR-specific formatting, we extracted transcribed words and speaker labels into a common format, consistent with what was used in Wang et al. 2024 [27]. These transcripts are referred to here as the ASR transcripts. The ASR transcripts were further standardized by removing punctuation and transforming all text to lowercase. 11% of reference transcripts were missing transcripts for part of the recording, ranging from 30 seconds up to 8 minutes and 27 seconds. For these, we trimmed the ASR transcripts to correspond with their reference transcripts.
While all input audio recordings contained conversation of 2 speakers, in some cases ASRs may only detect 1 speaker. This can be due to poor audio quality, noise in the recording, volume of the speakers or other factors. Any ASR transcripts with only one or more than two detected speakers were removed from the dataset. Some ASR transcripts, particularly those from WhisperX, had continuously repeating word sequences [41]. For example, words such as ‘and’ or phrases such as ‘that’s right’ would repeat consecutively such that the sentence would lose meaning. To minimize the effect of repeating sequences on model fine-tuning and testing, transcripts that contained more than 10 repetitions of a sequence were removed.
To create the training data, a transcript-preserving speaker transfer algorithm (TPST) [27] was used to transfer speaker labels from the reference transcripts to the ASR transcripts shown in Figure 3. The TPST algorithm accepts speaker-labeled source text and speaker-labeled target text. It aligns the two so that the source text speaker labels match the target text. The resulting transcripts are referred to here as the oracle transcripts. Oracle transcripts have the same text as the ASR transcripts but the speaker labels of the reference transcripts. They allow for comparison of diarization accuracy between an ASR transcripts and the ground-truth speaker labeling in the reference transcripts by removing effects for differences in the transcription itself.
@@ 2.4.2 ASR-specific models
We fine-tuned each of the three ASR-specific models using Amazon SageMaker on an ml.g5.8xlarge instance (1 NVIDIA A10G Tensor Core GPU with 24 GiB of memory) for 2 epochs, with a batch size of 6 and 5 gradient accumulation steps. We used Quantized Low-Ranked Adaption of Language Models [43] and Flash attention [44] for efficient fine-tuning, balancing computational resources with model performance improvement. The three resulting fine-tuned models are referred to here as the AWS model, Azure model, and WhisperX model, or collectively as the ASR-specific models.
@@ 2.5 Completion parser
A completion parser was developed to post-process the model output. It extracts only the relevant part from the model output, removing pre- and post-fixes if present, extracts speaker labels from the model output, and transfers those labels back to input text. Since LLMs can hallucinate, it is common that the LLM may alter original wording in the transcript in addition to speaker labeling. To mitigate this, the TPST algorithm––the same algorithm as seen in creating the oracle transcripts during training––was used to ensure speaker labels outputted by the model matched the original wording. This is done by transferring the completion speaker labels to the input ASR transcription wording, thus mitigating word changes made by the LLM in the completion wording. This completion parser was integrated as part of the model’s methods to ensure no word substitutions or modifications were made through use of the model.
@@ 2.6.1 Measurement of accuracy
Diarization accuracy was measured using two established metrics: delta concatenated minimum-permutation word error rate (deltaCP) [47] and delta speaker-attributed word error rate (deltaSA) [48]. For each, word error rate (WER) first needed to be calculated. WER was measured as the percentage of transcription errors divided by the total number of words. When measuring accuracy in this manuscript, deltaCP and deltaSA were calculated with the reference transcripts serving as ground truth. Together, deltaCP and deltaSA quantify how much additional error is introduced by speaker labeling, allowing us to differentiate between transcription errors and those stemming from misattributed speakers.
To calculate deltaCP, concatenated minimum-permutation word error rate (cpWER) was first calculated. To do this, each speaker’s words were separated into their own transcripts. For each possible permutation of the speaker’s words, the WER was calculated against the reference speaker transcript. The smallest observed WER for each of the two speakers was averaged. cpWER is considered the minimum possible WER among all permutations of the two speakers. deltaCP is simply cpWER subtracted by the original WER (i.e. deltaCP = cpWER - WER). The difference allows us to analyze errors introduced by speaker labeling, independent of the original WER.
To calculate deltaSA, speaker-attributed word error rate (SA-WER) was first calculated. To do this, each speaker’s words were separated into their own transcripts. WER was calculated against the reference speaker-transcripts. The WER of each of the two speakers was averaged. deltaSA is simply SA-WER subtracted by the WER (i.e. deltaSA = SA-WER - WER).
@@ 2.6.2 Assessing performance
Model performance was assessed by calculating deltaCP and deltaSA before and after diarization correction. Seven different models for diarization correction were used:
@@
DiarizationLM 8b Fisher v2 is the best performing model from the models that are made openly available, proposed by Wang et al. 2024 [27]. The framework they employed is the one we based our training strategy on for our ASR-specific models. The base model used was Llama3 with 8b parameters which is similar in size with the base model we used which had 7b parameters.
@@
In addition, when evaluating DiarizationLM 8b Fisher v2 which was proposed and open-sourced in Wang et al. 2024 [27], the following prompt template was used (as shown at https://huggingface.co/google/DiarizationLM-8b-Fisher-v2):
Prompt: [ASR transcript] $\rightarrow$
@@ 3.2 Word error rate
We looked at WER across the three ASRs. The WER across ASRs is shown in Table 1. The PriMock57 dataset had considerably better audio quality, given the mock sessions were carefully conducted in an isolated environment, which explains the improved transcription WER compared to the Fisher test set. Azure’s Speech to Text model performed best in terms of transcription accuracy in the Fisher test set, while AWS’s Transcribe performed the best in the PriMock57 dataset. WhisperX performed the worst in both datasets.
@@ 3.3 Baseline diarization accuracy
Diarization accuracy was compared between ASR transcripts and corresponding reference transcripts without any diarization correction. The baseline diarization accuracy across ASRs is shown in Table 1. In the Fisher dataset, AWS performed best, while Azure was next and WhisperX performed worst. On the PriMock57 dataset Azure performed the best and AWS was next. WhisperX performed worst on the PriMock57 dataset as well. The difference in diarization performance between AWS’s Transcribe and Azure’s Speech to Text is indicative of the fact that different ASRs may work better for different use cases.
Table 1: Baseline transcription and diarization accuracy measures across ASRs.
AWS transcripts Azure transcripts WhisperX transcripts
deltaCP deltaSA WER deltaCP deltaSA WER deltaCP deltaSA WER
Fisher test set 0.93 2.5 22.04 1.91 3.06 16.99 4.46 5.77 22.39
PriMock57 dataset 2.18 3.56 12.43 0.82 1.97 14.28 4.96 6.56 18.69
@@ 3.5 Fine-tuned model performance
The performance of the ASR-specific models and the ensemble model in improving diarization accuracy was measured against baseline in the Fisher test set. We also evaluated DiarizationLM 8b Fisher v2 [27]. The results are shown in Table 3.
@@
DiarizationLM 8b Fisher v2 is unable to improve diarization performance in unseen ASRs. Fine-tuned models markedly improved diarization accuracy compared to baseline in most cases.
@@
The ensemble model achieved the best overall performance. In the AWS transcripts, it had marginally worse performance compared to the AWS model. However, in the Azure and WhisperX transcripts, it performed better than even the ASR-specific models for Azure and WhisperX. Ablation studies on ensembles of two ASR-specific models indicate that while partial improvements can be achieved by combining subsets of ASR-specific models, the full three-model ensemble provides the best generalization across different ASR outputs (see C for details).
@@ 4 Discussion
In examining the results reported in Wang et al. 2024 [27] (using USM + turn-to-diarize as their ASR), their best performing model (PaLM 2-S finetuned (hyp2ora)) achieves an impressive 75% improvement in deltaCP, which significantly outperforms our ASR-specific models. However the best performing open-source model (Llama 3 8B finetuned (mixed) v2; named DiarizationLM 8b Fisher v2 in HuggingFace) achieves approximately 49% improvement in deltaCP, which is more in line with the performance gains seen in our AWS and Azure models when evaluated on their own test sets, that show 46% and 54% improvements respectively.
However, when evaluating DiarizationLM 8b Fisher v2 on transcriptions from unseen ASRs (using our test sets transcribed with AWS, Azure, and WhisperX) the model fails to generalize, leading to a deterioration in diarization performance. This decline in accuracy may be attributed to overfitting to source ASR, large differences in ASR transcription or differences in the amount of diarization errors. Notably, the WhisperX model, which shows the least degradation in performance, had the most comparable baseline diarization error (deltaCP = 4.46) to the ASR used in Wang et al. (deltaCP = 5.71), suggesting that diarization similarity plays a key role in the model’s ability to maintain performance.
@@
We acknowledge certain limitations, opening up opportunities for future work. First, a more thorough analysis of the diarization errors produced by each ASR would clarify the need for generalized approaches and help us understand better the types of errors each LLM is more probable to correct. In particular, future work should explore how the results of different ASR models are similar or different, and how LLMs use these variations to improve diarization performance. Especially as it becomes easier to quantify these errors, investigating these differences could provide deeper insights into model adaptability and generalization across ASR outputs. Second, the project was limited to transcripts in English. This restricts diarization correction to a single language. Future work could test this model on transcripts in different languages and also include additional languages in the dataset used for fine-tuning. Third, our experimentation with varied prompts was not extensive [53]. Given we were building on work from Wang et al. 2024 [27], we used the same prompts utilized in their work. Future work exploring the transferability of the model could include contextual information (e.g., this is a conversation between a doctor and a patient; this is a phone conversation between a salesperson and a customer) [54]. Finally, though our project focused on such models serving as post-processing steps, integrating multimodal data will allow for more robust diarization systems. Combining acoustic information with semantic insights from the LLM will enhance the system’s ability to label speech more effectively [31], especially in complex and noisy environments.
@@ 5 Conclusion
We investigated the use of fine-tuned LLMs to improve speaker diarization in conversational transcripts with existing speaker labels. Our findings demonstrate that LLMs can markedly improve diarization accuracy when specifically trained to do so. However, if such training is specific to one ASR tool or using a single dataset, it may constrain the model’s performance to transcripts only from that ASR tool and with similar diarization error rates. An ensemble model with knowledge across ASR tools allows for a generalizable and ASR-agnostic diarization correction tool. We believe such a model can be a useful post-processing step in applications that depend on transcription and diarization of conversations with multiple speakers. We have made these models publicly available at HuggingFace at https://huggingface.co/bklynhlth and built a Python function for easier use of the model available at https://github.com/bklynhlth/openwillis.
@@ 6 Acknowledgements
Though referenced throughout the manuscript, the authors would like to acknowledge the authors of Wang et al. 2024 [27], whose work formed the foundation of this manuscript. Much of the code used during this work, including the TPST algorithm, critical to model fine-tuning and testing, was originally developed by their group and kindly shared with us for replication.
@@ Appendix B Generalization to unseen ASR
Appendix C Ensemble model - Ablation study
The ensemble model was constructed by combining all three ASR-specific models (AWS, Azure, and WhisperX). Each individual model was fine-tuned on transcripts generated by a specific ASR, making them specialized for correcting diarization errors inherent to that ASR. The ensemble approach aimed to leverage the strengths of all three models to improve generalization across different ASRs.
To further analyze the contribution of each ASR-specific model to the ensemble’s performance, we conducted an ablation study evaluating ensembles composed of only two ASR-specific models at a time. The following combinations were tested:
  • 1.
    AWS + Azure
  • 2.
    AWS + WhisperX
  • 3.
    Azure + WhisperX
Each of these two-model ensembles was evaluated on the Fisher test set from all three ASRs to determine their effectiveness compared to the ASR-specific models and the full three-model ensemble. The results are shown in Table 7.
Table 7: Evaluation of ASR-specific and ensemble models on correction of diarization errors in the Fisher test set. Two-ASR ensemble models showed partial improvements but lacked full generalization. The full ensemble model, incorporating all three ASR-specific models, achieved the best overall performance across different ASRs, demonstrating improved robustness in diarization correction.
AWS transcripts Azure transcripts WhisperX transcripts
deltaCP deltaSA deltaCP deltaSA deltaCP deltaSA
Baseline 0.93 2.5 1.91 3.06 4.46 5.77
AWS model 0.5 -46% 2.05 -18% 1.3 -32% 2.41 -22% 4.11 -8% 5.39 -7%
Azure model 1.04 +12% 2.6 +4% 0.87 -54% 1.92 -37% 3.89 -13% 5.12 -11%
AWS+Azure model 0.54 -42% 2.04 -18% 0.91 -52% 1.99 -35% 3.45 -23% 4.64 -20%
AWS+WhisperX model 0.55 -41% 2.01 -20% 0.97 -49% 2.05 -33% 3.01 -33% 4.17 -28%
Azure+WhisperX model 1 +8% 2.54 +2% 0.77 -60% 1.81 -41% 2.95 -34% 4.08 -29%
The two-ASR ensembles performed better in the datasets which were included in their training sets compared to the third left out ASR, in some cases even surpassing the full ensemble model. This further underscores the challenge of achieving broad generalizability. Notably, the Azure+WhisperX ensemble performed the best in both Azure and WhisperX transcripts but failed to generalize to AWS transcripts, where it not only underperformed but also degraded baseline performance. This suggests that AWS transcripts exhibit distinctive characteristics compared to Azure and WhisperX transcripts, which appear to be more similar to each other. These structural differences likely contributed to the observed performance degradations.
Overall, the full ensemble model (AWS + Azure + WhisperX) provided a more balanced approach to diarization correction, reducing overfitting to any single ASR’s error patterns and ensuring better adaptability across different ASR outputs. Interestingly, the AWS+WhisperX ensemble demonstrated comparable or even superior performance in most cases, suggesting that the significant disparity between AWS and WhisperX transcripts in baseline diarization performance (deltaCP: 0.93 vs. 4.46; deltaSA: 2.5 vs. 5.77) may have encouraged complementary error correction.
These findings highlight the importance of incorporating diverse ASR-specific models into the ensemble that can compensate for each other’s weaknesses. While two-ASR ensembles offered performance gains over individual models, they lacked the full generalization benefits observed in the three-model ensemble. Future work could explore alternative fusion methods and incorporating more ASR-specific models to further enhance diarization correction in unseen ASR scenarios